Phil Dunphy, a real estate agent, is considering whether he should list an unusual $760,463 house for sale. If he lists it, he will need to spend $5,132 in advertising, staging, and fresh cookies. The current owner has given Phil 6 months to sell the house. If he sells it, he will receive a commission of $24,340. If he is unable to sell the house, he will lose the listing and his expenses. Phil estimates the probability of selling this house in 6 months to be 73%. What is the expected profit on this listing?

Answer by Theo(13342)

probability of selling the house is .73
his expected profit will be .73 * 24,340 - 5,132 = 12,636.2
the rationale is as follows:
if he tried to sell a house that cost as much as the one he is now trying to sell 100 times, he would sell the house 73 times and he would not sell the house 27 times.
each time he tried, it would cost him 5,132.
he would make 73 * 24,340 = 1,776,820 and he would lose 100 * 5,132 = 513,200.
he would net the difference between what he would make minus the cost.
that would be equal to a net of 1,263,620.
that would be an average of that / 100 = 12,636.2 per attempt.